OpenStudy (anonymous):

how do you solve sec^2x=tanx+1

OpenStudy (anonymous):

1st step...replace sec^2(x) with tan^2(x) + 1

OpenStudy (anonymous):

Okkk so here is how I wld solve this: we know that \( \sec^2x= \large \frac{1}{\cos^2x}\)

OpenStudy (anonymous):

and \( \tan x = \large \frac{ \sin x}{ \cos x}\) so \[ \sec^2x= \tan x +1 \] \[\frac{1}{ \cos ^2x}=\frac{ \sin x}{ \cos x}+1 \] Now we will multiply both sides of the equation by \( \cos ^2 x \) and we get: \[1= \sin x \cos x +\cos^2x\] One of the trig identities is \( 1= \sin^2x+ \cos^2x\) Therefore we replace the 1 with the identity and get: \[\sin^2x +\cos^2x=\sin x \cos x +\cos^2x\] We subtract \( \cos^2 x\) from both sides and get: \[\sin^2x=\sin x +\cos x \] Then we divide both sides by \( \sin x\) and \[ \sin x=\cos x\] Then we divide both sides by \( \cos x \) and finally get \[ \frac{ \sin x}{ \cos x} =1\] Which is \( \tan x =1 \) Then we multiply by sides by inverse tan and get: \[ \tan ^{-1} x=1\] Ummm and as you know at \( \tan (45)=1\) So our x=45 degrees

OpenStudy (anonymous):

Do you follow?

OpenStudy (anonymous):

thanks BlackLabel

OpenStudy (anonymous):

Ummm do you follow what i did? Any questions?